# The Dynamical History of the Kepler-221 Planet System

### v1

@@ 3 Charting the Course: A Successful Simulation
In this section, a particular simulation has been selected to demonstrate the four-phase model. The successful result in Figure 2 is not guaranteed and the success rate of each phase is discussed in detail in Section 4.4. Obviously, the successful outcome is somewhat tuned and parameter-dependent. Specifically, several key milestones can be identified. These include: the high eccentricity the planets acquire post-disk phase; the occurrence of a collision between planets d1 and d2; the reformation of the b/c two-body resonance, followed by the 3BR between b/c/e; the ability for sustained orbital expansion of the b/c/e system towards the current period ratios; and the inability of planet to dislodge these planets out of the 3BR they are in today. However, the successful result in Figure 2 serves the purpose of charting the course in this section. In the following, we will discuss in detail the physical conditions that must be satisfied to pass each successive milestone and quantify the overall success rate of each model step.
@@ 4 Early Resonance Formation, Dynamical Instability, and 3BR Reformation (Phase I-III)
In the previous section, we introduced a consistent simulation throughout all phases which successfully reproduced the observed configuration. In this section, the evolution from Phase I to III until the b/c/e resonance reformation will be investigated. The masses of planets in the simulations follow Model M2a in Table 2 in this section. The focus of this section will be on the conditions for which planet $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ manage to merge and the b, c, and e 3BR to successfully reform after the assumed planetary collision in Phase II. The range of parameters for the simulation from Phase I to III are listed in Table 3.
@@ 4.1 Evolution until Dynamical Instability
In the disk phase, the five planets (planets b, c, $\mathrm{d}_{1}$, $\mathrm{d}_{2}$, and e) form a stable first-order resonance chain by disk migration with planets b, c, and e in 6:3:1 resonance, as shown in panels a and e in Figure 2. The mass of planet d is split between $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ such that its center of mass lies close to the observed position of planet d after the orbital expansion phase. Therefore the mass ratio of $m_{\mathrm{\mathrm{d}_{1}}}/m_{\mathrm{\mathrm{d}_{2}}}$ is assumed to be in the range of $[1,1.04]$. The formation of the 1st-order resonance chain is guaranteed because the planets are initialized close to their corresponding resonance position. To test the stability of the resonance chain after resonance formation, we apply a disk dispersal timescale $t_{\mathrm{d}}$ in the range of $[50\,\mathrm{yr},10^{4}\,\mathrm{yr}]$. The planets remain in resonance over the duration of disk dispersal. This is because the number of planets in the system is lower than the critical number in a 1st-order resonance chain (Matsumoto & Ogihara, 2020). Therefore, dynamical instability is required to break the resonance chain.
@@
It has been suggested that dynamical instability (Izidoro et al., 2017) or planetesimal scattering (Raymond et al., 2022) could perturb the system stability and break the resonance chain. Here, we do not model the specifics of the resonance breaking process, but simply enforce it by perturbing planet $\mathrm{d_{2}}$ with a ”kick” just strong enough to break the $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ resonance. In the simulation, we apply a small instantaneous torque on planet $\mathrm{d_{2}}$ which decreases its semi-major axis ($\Delta a/a$) in the range of $[0.5\%,1.2\%]$ within $10^{-3}\mathrm{yr}$. The simulations show a $\Delta a/a$ as small as $0.7\%$ breaks the entire resonance chain (Zhou et al., 2007; Pichierri & Morbidelli, 2020) and results in the rapid merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$, as shown in panels b and f in Figure 2.
@@
In the collision phase, a quick merging between planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ is preferred. In that case, planets b/c/e would not experience much perturbation of their semi-major axes and would stay close to the 3BR resonance position, which is beneficial for reformation (see below). On the other hand, if the collision phase is long, the planets would undergo a long time of instability that would cause the positions of planets b/c/e to deviate significantly from the exact integer period ratio. This is disadvantageous for the subsequent 3BR reformation process. How rapid $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ merge after resonance breaking is determined by the eccentricity of planets ${\mathrm{d}_{1}}$ and ${\mathrm{d}_{2}}$. If the initial eccentricity is high (e.g., around 0.1), orbit crossing and merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ will quickly follow the breaking of resonance. On the other hand, if the eccentricity remains low, the orbit of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ would only cross after the planet eccentricity excites and the time required for the planet merging will be much longer (Tamayo et al., 2021). Also, a higher eccentricities of planets in the resonance chains implies that they are deeper in resonance with the period ratio closer to integer (Huang & Ormel, 2023). This is also advantageous for the 3BR reformation afterward with the initial position of the planet closer to the b/c/e 3BR position. The ratio between the semi-major axis damping timescale and eccentricity damping timescale $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}$ determines the eccentricity of the planets after disk dispersal (Teyssandier & Terquem, 2014; Papaloizou et al., 2018) with the eccentricity being proportional to the square-root of the damping ratio parameter. For example, in 45% of the simulations planets d1 and d2 would merge within $10^{3}\,\mathrm{yr}$ with $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}=250$ while the merging rate within $10^{3}\,\mathrm{yr}$ is lower at 10% when $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}=450$. Therefore, in 70% of the simulations with $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}=450$, planets b/c/e exhibit a large deviation from 3BR with $B_{\mathrm{norm}}>0.015$ after the merging of planets d1 and d2 (see Equation (4)), making the 3BR reformation afterward almost impossible.
@@
Figure 3: Success and failure in reforming the b, c, and e 3BR post-collision. Top panels show the evolution of $P_{\mathrm{c}}/P_{\mathrm{b}}$ and $P_{\mathrm{e}}/P_{\mathrm{c}}$. The black dashed line and red dashed line correspond to the 0th-order and 1st-order 3BR, respectively, and the color of the dots indicates time. The bottom panels show the b/c/e 3BR angle. In panels a and d, there is no torque on planet b, and $Q_{\mathrm{phy}}$ is the same for all planets. The planets cross the 0th-order 3BR and get trapped in a 1st-order 3BR, which is inconsistent with the present state. In panels b and e, planet b experienced an exponentially decaying outward torque with a timescale of $5\,\mathrm{Myr}$ with the same $Q_{\mathrm{phy}}$ for all planets. In panels c and f, tidal dissipation is more sufficient in planet c with $Q_{\mathrm{c,phy}}/Q_{\mathrm{b,phy}}=0.1$.
@@
The success rate of mass model M2a and M2d (see Table 2 is tested in the simulation with the parameters in the range of Table 3. With mass model M2a, in 51% of the simulations planets d1 and d2 merge within $10^{4}\mathrm{yr}$. In the other 49% of the simulation, planets b/c/e already deviate a lot from the integer period ratio, making the future 3BR reformation impossible, so the simulation is truncated. On the other hand, the merging of other planets is possible when mass model M2d is assumed because of more massive planets d1 and d2. In 47% of the simulations still d1 and d2 merge within $10^{4}\,\mathrm{yr}$. In 6% of the simulations planets d2 and e merge while in 11% of the simulation planets c and d1 merge. The other simulations are truncated since planets b/c/e already deviate from 3BR position.
@@ 4.2 Resonance Reformation after Dynamical Instability
Panels c and f of Figure 2 represent the post-collision phase after the merging of planets d1 and d2. While the success rate in Phase I and II is relatively high, the reformation of the b/c/e 3BR after dynamical instability is much more challenging. Figure 2 features in fact a successful condition in which the b, c, and e 3BR reforms because of an outward (positive) torque applied on planet b. This ensures that planets b and c convergently approach each other, facilitating 2BR formation (see Section 4.3). Otherwise, the reformation of the (2, 3, 5) 0th-order 3BR cannot be accomplished (Petit, 2021). Specifically, the desired scenario is that planets b and c first reform the 2:1 two-body resonance under convergent migration, which is relatively straightforward. Then b, c, and e 3BR form at the same time with the c and e in 3:1 resonance. However, in the scenario where $Q_{\mathrm{phy}}$ is the same for all planets, planets b and c tend to experience divergent migration. As a result, there is only a slight chance to reform the b, c, and e 3BR.
We run over 50 simulations with the setup applying the same $Q_{\mathrm{phy}}$ (see Section 2.3) on all planets but with slightly different initial positions and eccentricities of the planets for each simulation. In all these simulations planets b, c, and e fail to reform the 3BR. An illustrative example is given by panels a and d in Figure 3. Instead of forming the desired 3BR, planet b/c/e either form a 1st-order 3BR (40%; as in Figure 3a) or the formation of the 3BR fails altogether (60%). The reason why a 1st-order 3BR forms but not the 0th-order is because the formation of a 1st-order 3BR is not preconditioned on the requirement of being close to exact resonance (see Appendix A) (Petit, 2021). Therefore, in the absence of convergent motions, the planets would either fail to be locked in resonance or form the 1st-order 3BR and expand along it (red dashed line in Figure 3).
@@ 4.3 Potential Mechanisms Facilitating Resonance Reformation
The first mechanism is a potential outward torque operating on planet b, which promotes the convergent migration between b and c and reforms the resonance. Panel b and e in Figure 3 show the 3BR reformation process corresponding to an exponentially decaying outward torque on planet b. The origin of the torque could include mass loss on planet b (Wang & Lin, 2023) or dynamical tides (Ahuir et al., 2021). In the simulation, we take the form of the exponentially decaying torque:
$\Gamma=\Gamma_{\mathrm{0}}\exp{\left(-\frac{t}{t_{\Gamma}}\right)}$ (8)
where $t$ is the simulation time after the merging of planets d1 and d2, $\Gamma_{0}$ is the value of the torque at $t=0$ and $t_{\Gamma}$ is the timescale on which the torque decays. The initial value of the torque can equivalently be parameterized in terms of an orbital decay timescale, e.g., if it acts on planet b:
$t_{a0}^{\mathrm{b}}\equiv\frac{L_{\mathrm{b}}}{2\Gamma_{0}}$ (9)
@@
Panel b of Figure 3 presents the period ratio diagram of a simulation with the imposed torque operating on planet b. It can be seen that the planet period ratio first oscillates around the resonance position relatively far from the integer period ratio. Then the positive torque on b decreases $P_{\mathrm{c}}/P_{\mathrm{b}}$ while quickly trapping planets b and c in two-body resonance. Planets b, c, and e also quickly get captured in 3BR, as can be seen from the librating resonance angle shown in panel e of Figure 3. Subsequently, the period ratio of the planets expands along the 0th-order resonance line.
The second mechanism to promote resonance reformation is more efficient damping on planet c. Planet c could potentially have a relatively small $Q_{\mathrm{phy}}$ (with $Q_{\mathrm{c,phy}}/Q_{\mathrm{b,phy}}=0.1$) due to efficient obliquity tides (Louden et al., 2023; Goldberg & Batygin, 2021) while the effective tidal $Q$ parameter for other planets is much larger (see Section 6.3). Tidal damping on planet c would result in its inward migration and contribute to the convergent migration between planets b and c, necessary to the formation of the b and c 2BR and the b, c, and e 3BR. Panel c and f in Figure 3 correspond to such a simulation. Here, damping on planet c is stronger, with a smaller tidal $Q$ parameter for planet c, ($Q_{\mathrm{c}}/Q_{\mathrm{b}}=0.1$). As shown in the figure, planets b, c, and e first oscillate around the 0th-order resonance (black dashed line) with large amplitude. Then tidal damping on the system migrates planets b and c inward, decreasing $P_{\mathrm{c}}/P_{\mathrm{b}}$ while increasing $P_{\mathrm{e}}/P_{\mathrm{c}}$. The subsequent convergent migration first forms the b and c 2BR. Afterward, tidal damping on planets b and c migrates c outward, leading to the convergent migration between planets c and e and the reformation of b/c/e 3BR.
The reformation of the b/c/e 3BR requires convergent migration between planets b and c. Otherwise, planets b, c, and e will be trapped in the 1st-order 3BR or will not get trapped in any 3BR. Here, we have demonstrated the viability of resonance reformation for two physically plausible scenarios of an outward torque on planet b and more effective tides on planet c. Nevertheless, the reformation of the b/c/e 3BR is stochastic and cannot be guaranteed. This will be discussed in more detail in Section 4.4.

### v2

@@ 3 Charting the course: a successful simulation
In this section, a particular simulation has been selected to demonstrate the four-phase model. The successful result in Figure 2 is not guaranteed and the success rate of each phase is discussed in detail in Section 4.4. Obviously, the successful outcome is somewhat tuned and parameter-dependent. Specifically, several key milestones can be identified. These include: the high eccentricity the planets acquire post-disk phase; the occurrence of a collision between planets d1 and d2; the reformation of the b/c two-body resonance, followed by the 3BR between (b, c, e); the ability for sustained orbital expansion of the (b, c, e) system toward the current period ratios; and the inability of planet to dislodge these planets out of the 3BR they are in today. However, the successful result in Figure 2 serves the purpose of charting the course in this section. In the following, we discuss in detail the physical conditions that must be satisfied to pass each successive milestone and quantify the overall success rate of each model step.
@@ 4 Early resonance formation, dynamical instability, and 3BR reformation (Phase I-III)
In the previous section, we introduced a consistent simulation throughout all phases which successfully reproduced the observed configuration. In this section, the evolution from Phase I to III until the (b, c, e) resonance reformation will be investigated. The masses of planets in the simulations follow Model M2a in Table 2 in this section. The focus of this section will be on the conditions for which planet $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ manage to merge and the b, c, and e 3BR to successfully reform after the assumed planetary collision in Phase II. The range of parameters for the simulation from Phase I to III are listed in Table 3.
@@ 4.1 Evolution until dynamical instability
In the disk phase, the five planets (planets b, c, $\mathrm{d}_{1}$, $\mathrm{d}_{2}$, and e) form a stable first-order resonance chain by disk migration with planets b, c, and e in 6:3:1 resonance, as shown in panels a and e in Figure 2. The mass of planet d is split between $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ such that its center of mass lies close to the observed position of planet d after the orbital expansion phase. Therefore the mass ratio of $m_{\mathrm{\mathrm{d}_{1}}}/m_{\mathrm{\mathrm{d}_{2}}}$ is assumed to be in the range of $[1,1.04]$. The formation of the first-order resonance chain is guaranteed because the planets are initialized close to their corresponding resonance position. To test the stability of the resonance chain after resonance formation, we apply a disk dispersal timescale $t_{\mathrm{d}}$ in the range of $[50\,\mathrm{yr},10^{4}\,\mathrm{yr}]$. The planets remain in resonance throughout disk dispersal. This is because the number of planets in the system is lower than the critical number in a first-order resonance chain (Matsumoto & Ogihara, 2020). Therefore, dynamical instability is required to break the resonance chain.
@@
It has been suggested that dynamical instability (Izidoro et al., 2017) or planetesimal scattering (Raymond et al., 2022) could perturb the system stability and break the resonance chain. Here, we do not model the specifics of the resonance breaking process, but simply enforce it by perturbing planet $\mathrm{d_{2}}$ with a ”kick” just strong enough to break the $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ resonance. In the simulation, we apply a small instantaneous torque on planet $\mathrm{d_{2}}$ that decreases its semi-major axis ($\Delta a/a$) in the range of $[0.5\%,1.2\%]$ within $10^{-3}\mathrm{yr}$. The simulations show a $\Delta a/a$ as small as $0.7\%$ breaks the entire resonance chain (Zhou et al., 2007; Pichierri & Morbidelli, 2020) and results in the rapid merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$, as shown in panels b and f in Figure 2.
@@
In the collision phase, a quick merging between planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ is preferred. In that case, planets (b, c, e) would not experience much perturbation of their semi-major axes and would stay close to the 3BR resonance position, which is beneficial for reformation (see below). On the other hand, if the collision phase is long, the planets would undergo a long time of instability that would cause the positions of planets (b, c, e) to deviate significantly from the exact integer period ratio. This is disadvantageous for the subsequent 3BR reformation process. How rapid $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ merge after resonance breaking is determined by the eccentricity of planets ${\mathrm{d}_{1}}$ and ${\mathrm{d}_{2}}$. If the initial eccentricity is high (for instance, around 0.1), orbit crossing and merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ will quickly follow the breaking of resonance. On the other hand, if the eccentricity remains low, the orbit of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ would only cross after the planet eccentricity excites and the time required for the planet merging will be much longer (Tamayo et al., 2021). Also, a higher eccentricity of planets in the resonance chains implies that they are deeper in resonance with the period ratio closer to integer (Huang & Ormel, 2023). This is also advantageous for the 3BR reformation afterward with the initial position of the planet closer to the (b, c, e) 3BR position. The ratio between the semi-major axis damping timescale and eccentricity damping timescale $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}$ determines the eccentricity of the planets after disk dispersal (Teyssandier & Terquem, 2014; Papaloizou et al., 2018) with the eccentricity being proportional to the square-root of the damping ratio parameter. For example, in 45% of the simulations planets d1 and d2 would merge within $10^{3}\,\mathrm{yr}$ with $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}=250$ while the merging rate within $10^{3}\,\mathrm{yr}$ is lower at 10% when $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}=450$. Therefore, in 70% of the simulations with $t_{\mathrm{a,i}}/t_{\mathrm{e,i}}=450$, planets (b, c, e) exhibit a large deviation from 3BR with $B_{\mathrm{norm}}>0.015$ after the merging of planets d1 and d2 (see Equation (4)), making the 3BR reformation afterward almost impossible.
@@
Figure 3: Success and failure in reforming the b, c, and e 3BR post-collision. Top panels show the evolution of $P_{\mathrm{c}}/P_{\mathrm{b}}$ and $P_{\mathrm{e}}/P_{\mathrm{c}}$. The black dashed line and red dashed line correspond to the zeroth-order and first-order 3BR, respectively, and the color of the dots indicates time. The bottom panels show the (b, c, e) 3BR angle. In panels a and d, there is no torque on planet b, and $Q_{\mathrm{phy}}$ is the same for all planets. The planets cross the zeroth-order 3BR and get trapped in a first-order 3BR, which is inconsistent with the present state. In panels b and e, planet b experienced an exponentially decaying outward torque with a timescale of $5\,\mathrm{Myr}$ with the same $Q_{\mathrm{phy}}$ for all planets. In panels c and f, tidal dissipation is more sufficient in planet c with $Q_{\mathrm{c,phy}}/Q_{\mathrm{b,phy}}=0.1$.
@@
The success rate of mass model M2a and M2d (see Table 2 is tested in the simulation with the parameters in the range of Table 3. With mass model M2a, in 51% of the simulations planets d1 and d2 merge within $10^{4}\mathrm{yr}$. In the other 49% of the simulation, planets (b, c, e) already deviate a lot from the integer period ratio, making the future 3BR reformation impossible, so the simulation is truncated. On the other hand, the merging of other planets is possible when mass model M2d is assumed because of more massive planets d1 and d2. In 47% of the simulations still d1 and d2 merge within $10^{4}\,\mathrm{yr}$. In 6% of the simulations planets d2 and e merge while in 11% of the simulation planets c and d1 merge. The other simulations are truncated since planets (b, c, e) already deviate from the 3BR position.
@@ 4.2 Resonance reformation after dynamical Instability
Panels c and f of Figure 2 represent the post-collision phase after the merging of planets d1 and d2. While the success rate in Phase I and II is relatively high, the reformation of the (b, c, e) 3BR after dynamical instability is much more challenging. Figure 2 features in fact a successful condition in which the b, c, and e 3BR reforms because of an outward (positive) torque applied on planet b. This ensures that planets b and c convergently approach each other, facilitating 2BR formation (see Section 4.3). Otherwise, the reformation of the (2, 3, 5) zeroth-order 3BR cannot be accomplished (Petit, 2021). Specifically, the desired scenario is that planets b and c first reform the 2:1 two-body resonance under convergent migration, which is relatively straightforward. Then b, c, and e 3BR form at the same time with the c and e in 3:1 resonance. However, in the scenario where $Q_{\mathrm{phy}}$ is the same for all planets, planets b and c tend to experience divergent migration. As a result, there is only a slight chance to reform the b, c, and e 3BR.
We run over 50 simulations with the setup applying the same $Q_{\mathrm{phy}}$ (see Section 2.3) on all planets but with slightly different initial positions and eccentricities of the planets for each simulation. In all these simulations, planets b, c, and e fail to reform the 3BR. An illustrative example is given by panels a and d in Figure 3. Instead of forming the desired 3BR, planet (b, c, e) either form a first-order 3BR (40%; as in Figure 3a) or the formation of the 3BR fails altogether (60%). The reason why a first-order 3BR forms but not the zeroth-order is because the formation of a first-order 3BR is not preconditioned on the requirement of being close to exact resonance (see Appendix A) (Petit, 2021). Therefore, in the absence of convergent motions, the planets would either fail to be locked in resonance or form the first-order 3BR and expand along it (red dashed line in Figure 3).
@@ 4.3 Potential mechanisms facilitating resonance reformation
The first mechanism is a potential outward torque operating on planet b, which promotes the convergent migration between b and c and reforms the resonance. Panels b and e in Figure 3 show the 3BR reformation process corresponding to an exponentially decaying outward torque on planet b. The origin of the torque could include mass loss on planet b (Wang & Lin, 2023) or dynamical tides (Ahuir et al., 2021). In the simulation, we take the form of the exponentially decaying torque:
$\Gamma=\Gamma_{\mathrm{0}}\exp{\left(-\frac{t}{t_{\Gamma}}\right)},$ (8)
where $t$ is the simulation time after the merging of planets d1 and d2, $\Gamma_{0}$ is the value of the torque at $t=0$ and $t_{\Gamma}$ is the timescale on which the torque decays. The initial value of the torque can equivalently be parameterized in terms of an orbital decay timescale, for instance, if it acts on planet b:
$t_{a0}^{\mathrm{b}}\equiv\frac{L_{\mathrm{b}}}{2\Gamma_{0}},$ (9)
@@
Panel b of Figure 3 presents the period ratio diagram of a simulation with the imposed torque operating on planet b. It can be seen that the planet period ratio first oscillates around the resonance position relatively far from the integer period ratio. Then the positive torque on b decreases $P_{\mathrm{c}}/P_{\mathrm{b}}$ while quickly trapping planets b and c in two-body resonance. Planets b, c, and e also quickly get captured in 3BR, as can be seen from the librating resonance angle shown in panel e of Figure 3. Subsequently, the period ratio of the planets expands along the zeroth-order resonance line.
The second mechanism to promote resonance reformation is more efficient damping on planet c. Planet c could potentially have a relatively small $Q_{\mathrm{phy}}$ (with $Q_{\mathrm{c,phy}}/Q_{\mathrm{b,phy}}=0.1$) due to efficient obliquity tides (Louden et al., 2023; Goldberg & Batygin, 2021) while the effective tidal $Q$ parameter for other planets is much larger (see Section 6.3). Tidal damping on planet c would result in its inward migration and contribute to the convergent migration between planets b and c, necessary to the formation of the b and c 2BR and the b, c, and e 3BR. Panel c and f in Figure 3 correspond to such a simulation. Here, damping on planet c is stronger, with a smaller tidal $Q$ parameter for planet c, ($Q_{\mathrm{c}}/Q_{\mathrm{b}}=0.1$). As shown in the figure, planets b, c, and e first oscillate around the zeroth-order resonance (black dashed line) with large amplitude. Then tidal damping on the system migrates planets b and c inward, decreasing $P_{\mathrm{c}}/P_{\mathrm{b}}$ while increasing $P_{\mathrm{e}}/P_{\mathrm{c}}$. The subsequent convergent migration first forms the b and c 2BR. Afterward, tidal damping on planets b and c migrates c outward, leading to the convergent migration between planets c and e and the reformation of (b, c, e) 3BR.
The reformation of the (b, c, e) 3BR requires convergent migration between planets b and c. Otherwise, planets b, c, and e will be trapped in the first-order 3BR or will not get trapped in any 3BR. Here, we have demonstrated the viability of resonance reformation for two physically plausible scenarios of an outward torque on planet b and more effective tides on planet c. Nevertheless, the reformation of the (b, c, e) 3BR is stochastic and cannot be guaranteed. This is discussed in more detail in Section 4.4.